OpenStudy (anonymous):

Geometric sequence: an=3(1/2)^(n-1) First term is 3 and common ratio is 1/2. Find sum of the first 3 terms, or S3. Either add the first 3 terms or use the formula: Sn=a1-a1^r^n/1-r. Help on how to do this?! Please!

OpenStudy (bahrom7893):

Use this formula: Sn=a1-a1^r^n/1-r, but just rewrite it what is that? Sn = (a1-a1^r)^n/(1-r) ?

OpenStudy (anonymous):

Yes, the one you typed is correct.

OpenStudy (bahrom7893):

a1 = 3 r = 1/2 n=3 S3 = ?

OpenStudy (bahrom7893):

Substitue all the values u have for Sn formula, wehre a1=3 r=1/2 n=3

OpenStudy (anonymous):

O.k, thank you so much :)